OpenStudy (anonymous):

HELP NEEDED PLEASE!!! Eliminate the parameter. x = 4 cos t, y = 4 sin t

OpenStudy (anonymous):

@mathstudent55 @SolomonZelman @zepdrix

OpenStudy (anonymous):

I'm kind of new to these also. Perhaps you could solve each for t and set them equal to one another?

OpenStudy (anonymous):

x=4cost => cost=x/4 and y=4sint => sint=y/4 now x^2/16+y^2/16 = cos^2t+sin^2t = 1

OpenStudy (anonymous):

So \[\cos^{-1}(\frac{x}{4}) = \sin^{-1}(\frac{y}{4})\]

OpenStudy (anonymous):

sorry: x^2/16+y^2/16+xy/8=1

OpenStudy (anonymous):

um..so \[\frac{ x ^{2} }{ 16}+\frac{ y^{2} }{ 16}+\frac{ xy }{ 8}=1\]

OpenStudy (anonymous):

I think I'm starting to fade. :P

OpenStudy (anonymous):

im just so confused

OpenStudy (anonymous):

Give me another minute, I think I'm almost there.

OpenStudy (vishweshshrimali5):

\[\large{\cos^2 t + \sin^2 t = 1}\]

OpenStudy (vishweshshrimali5):

\[\large{x = 4\cos t}\] \[\large{\implies \cos t = \cfrac{x}{4}}\] Similarly, \[\large{\sin t = \cfrac{y}{4}}\]

OpenStudy (anonymous):

Ah duh! Thanks vishweshrimali5! That's it. :)

OpenStudy (vishweshshrimali5):

Substitute these values of cos t and sin t in : \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 \[\large{\cos^2 t + \sin^2 t = 1}\] \(\color{blue}{\text{End of Quote}}\) And you will get your answer :)

OpenStudy (anonymous):

thank you :)

OpenStudy (vishweshshrimali5):

Your welcome :)